hello again i have a problem i need help with. the sum of two numbers is 89 adding four times the first number to the second number gives 155 which one of the equations expresses this problem algebraically? A. (X-89)+4X=155 B. 89+4x=155 C.4x+(89-x)=155 D.4(89+x)+x=155 i picked B thank you for all of your help : hello again i have a problem i need help with. the sum of two numbers is 89 adding four times the first number to the second number gives 155 which one of the equations expresses this problem algebraically? A. (X-89)+4X=155 B. 89+4x=155 C.4x+(89-x)=155 D.4(89+x)+x=155 i picked B thank you for all of your help
Answer by Fombitz(1270)

It would be more helpful if you wrote down the steps that got you to your answer.
Anyways, let's go through it step by step.
First name your unknown numbers.
We'll use x and y.
"The sum of two numbers is 89."
1.x+y=89
adding four times the first number to the second number gives 155
2.4*x+y=155
Now you have two equations.
You can use eq. 1 to substitute for y in eq. 2,
1.x+y=89
y=89-x
Now substitute into eq. 2,
2.4*x+y=155
4*x+(89-x)=155
This line looks like C.